Figure 1: Antilock brake system components. 1 Figure 2: Road coefficient of adhesion versus wheel slip curves for different road surfaces.

2 Figure 3: Front wheel free body diagram; for rear wheel F_{bf} becomes F_{br}.

3 Figure 4: ABS operating modes.

4